OpenStudy (anonymous):

can someone help me understand how to do this I have 3 then a radical -216

OpenStudy (anonymous):

\[\sqrt[3]{-216}\]

OpenStudy (mertsj):

That 3 is the index. It means to take the 3rd root of -216. It means, "What number makes this sentence true": \[(?)^3=-216\]

OpenStudy (whpalmer4):

The value of that expression is some number x, such that x*x*x = -216.

OpenStudy (mertsj):

Because taking the third root is the inverse of raising to the third power.

OpenStudy (anonymous):

oooo, so can I use prime factorization for this?

OpenStudy (whpalmer4):

Here's a hint: \[\sqrt[3]{-8} = -2\]

OpenStudy (mertsj):

That might be helpful.

OpenStudy (anonymous):

Am I suppose to write and guess until I get something that hits 216? Because this seems ineffective

OpenStudy (mertsj):

Yes, sort of. First of all, we know it has to be an even number because an odd number raised to a power could not be even. Secondly, we know it has to be negative.

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

So, pick up your calculator and start with -2 and raise it to the third power. Do you get -216?

OpenStudy (anonymous):

no I dont want to get in the habit of using a calculator and the teacher doesnt like it either

OpenStudy (mertsj):

Ok. Now try -4. What is -4 times -4 times -4?

OpenStudy (whpalmer4):

You could also factor 216: 216=2*108 = 2*2*54 = 2*2*2*27 = 2*2*2*3*9 = 2*2*2*3*3*3

OpenStudy (anonymous):

how does that help me find out what number cubed is 216

OpenStudy (whpalmer4):

Well, for every group of 3 identical factors, you keep one. Multiply what's left together. For example, 27 = 3*9 = 3*3*3. We've got 3 3's, so we keep one. Well, 3*3*3 = 27, so 3^3 = 27.

OpenStudy (anonymous):

ok I get that part

OpenStudy (whpalmer4):

How about finding the cube root of 64? 64 = 2*2*2*2*2*2. We can write that as (2*2*2)*(2*2*2) and if we keep 1 out of every 3, that means 2*2 = 4 is the cube root of 64.

OpenStudy (whpalmer4):

How about 1000? 1000 = 2*500 = 2*2*250 = 2*2*2*125 = 2*2*2*5*25 = 2*2*2*5*5*5. You've got 3 each of 2 and 5. Keep 1 of each, multiply them together, 2*5 = 10. 10^3 = 1000. 10 is the cube root of 1000. Get the picture?

OpenStudy (whpalmer4):

Handy rules for determining divisibility when factoring: if the number is even, it is divisible by 2 if the number ends in 5 or 0, it is divisible by 5 if the sum of the digits is a multiple of 3, it is divisible by 3

OpenStudy (anonymous):

hmm ok I think I get what your saying I got a little lost when u got to the higher numbers but thats cause my mind just shut down let me look back at what you said

OpenStudy (anonymous):

oooo ok I think I got it.

OpenStudy (whpalmer4):

It's just like for square root, except instead of needing pairs of identical factors, you need triples...

OpenStudy (anonymous):

so let me try another one

OpenStudy (whpalmer4):

How about 512?

OpenStudy (anonymous):

u want the cube index? or squared

OpenStudy (whpalmer4):

cube root of 512, since that's what you were doing in this problem.

OpenStudy (anonymous):

ooo I got stuck because I was factoring and I thought it had to factor evenly but I got it so 8 cubed is 512

OpenStudy (whpalmer4):

Correct!

OpenStudy (anonymous):

hmmm ok so I was trying for 5 radical -243 and I got stuck look at what I did